# Explicit Abstention Knobs for Predictable Reliability in Video Question Answering

### v1

@@ Appendix D Prompt Template
System Prompt (v1) You are answering a multiple-choice video question using ONLY the provided frames. Rules: Use only visual evidence from the frames. Do not assume details not visible. Do not explain your reasoning. Do not write prose. If the frames do not contain enough evidence to choose confidently, abstain. If you do not abstain, you must select exactly one option A–E. Your confidence must be a number in $[0,1]$ reflecting probability of correctness. Output Format: JSON only with this exact schema (no extra keys): 1{ 2 "choice": "A" | "B" | "C" | "D" | "E" | null, 3 "confidence": <number>, 4 "abstain": <boolean>, 5 "evidence_span": [start_idx, end_idx] | null 6}

### v2

@@ E.1 Self-Reported Confidence Prompt
JSON Prompt for Self-Reported Confidence (v1) You are answering a multiple-choice video question using ONLY the provided frames. Rules: Use only visual evidence from the frames. Do not assume details not visible. Do not explain your reasoning. Do not write prose. If the frames do not contain enough evidence to choose confidently, abstain. If you do not abstain, you must select exactly one option A–E. Your confidence must be a number in $[0,1]$ reflecting probability of correctness. Output Format: JSON only with this exact schema (no extra keys): 1{ 2 "choice": "A" | "B" | "C" | "D" | "E" | null, 3 "confidence": <number>, 4 "abstain": <boolean>, 5 "evidence_span": [start_idx, end_idx] | null 6}
E.2 Logprob Extraction Prompt
Letter-Only Prompt for Logprob Extraction You are answering a multiple choice question about a video. The video frames are provided in chronological order. Based on the video frames, answer the following question by selecting exactly ONE option (A, B, C, D, or E). Question: {QUESTION_TEXT} Options:
A) {OPT_A}
B) {OPT_B}
C) {OPT_C}
D) {OPT_D}
E) {OPT_E}
Respond with ONLY a single letter (A, B, C, D, or E) and nothing else.
The letter-only prompt is used with response_logprobs=True to extract token-level probabilities over answer options. This simpler output format avoids JSON parsing and enables direct access to the model’s softmax distribution over the five answer tokens.
E.3 Logprob-to-Probability Computation
Figure 9 illustrates the logprob extraction pipeline. Raw log probabilities $\ell_{i}$ from the model’s token distribution are normalized via softmax over the five answer options A–E. This renormalization is necessary because the model’s full vocabulary distribution includes tokens beyond the answer options; we restrict to the answer space to compute a proper probability distribution. The resulting $p_{i}$ values sum to 1 and represent the model’s relative preference among answer choices.
Figure 9: Logprob-to-probability computation pipeline. Video frames and prompt are processed by the transformer model. The final hidden state is projected via the LM head to produce logits over the vocabulary. We extract logprobs for tokens A–E and renormalize via softmax to obtain a probability distribution over answer options, from which we derive $p_{\max}$, margin, and entropy.